# Scaling Laws for Reward Model Overoptimization in Direct Alignment Algorithms

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has been crucial to the recent success of Large Language Models (LLMs), however, it is often a complex and brittle process. In the classical RLHF framework, a reward model is first trained to represent human preferences, which is in turn used by an online reinforcement learning (RL) algorithm to optimize the LLM. A prominent issue with such methods is reward over-optimization or reward hacking, where performance as measured by the learned proxy reward model increases, but true quality plateaus or even deteriorates. Direct Alignment Algorithms (DDAs) like Direct Preference Optimization have emerged as alternatives to the classical RLHF pipeline by circumventing the reward modeling phase. However, although DAAs do not use a separate proxy reward model, they still commonly deteriorate from over-optimization. While the so-called reward hacking phenomenon is not well-defined for DAAs, we still uncover similar trends: at higher KL budgets, DAA algorithms exhibit similar degradation patterns to their classic RLHF counterparts. In particular, we find that DAA methods deteriorate not only across a wide range of KL budgets but also often before even a single epoch of the dataset is completed. Through extensive empirical experimentation, this work formulates and formalizes the reward over-optimization or hacking problem for DAAs and explores its consequences across objectives, training regimes, and model scales.
@@ 3 Empirical Analysis of Overoptimization in DAAs
First, we examine the over-optimization problem in DAAs and compare it to those observed in traditional RLHF methods. All our experiments are carried out using the Reddit TL;DR summarization dataset [55] and the Pythia family of Large Language Models [6].
@@ 4 Reward Exploitation in Direct Alignment Algorithms
While the phenomena observed in the previous section echo those observed in classical RLHF, their underlying causes may be distinct. Reward over-optimization in classical RLHF is largely attributed to querying a proxy reward function that is potentially OOD, while DAAs do not train a separate reward model. Instead, DAAs are generally understood as fitting an “implicit” reward model to preference data with the parameterization $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}{(y|x)}}$ using the objective in eq. 2. Therefore, the OOD behavior of the policy is inextricably linked to the OOD behavior of the implicit reward model. Here we demonstrate that the reward modeling objective used is heavily under-constrained, allowing for a potentially large number of solutions that can place weight on OOD responses. This is especially problematic for DAAs which deterministically map the optimal policy from the “implicit” reward.
@@
Rank Deficiency with Finite Preferences. In DAAs, the language modeling problem is treated as a contextual bandit. However, the space of possible prompts $x\in\mathcal{X}$ and answers $y\in\mathcal{Y}$ are both exponentially large in sequence length. However, as highlighted by Tang et al. [59], DAAs often assume full support of the reference distribution when mapping from the implicit reward to the optimal policy $\pi$ by eq. 10. However, in practice such coverage is impossible. Instead, preference datasets cover a minuscule portion of the prompt-response space. Unfortunately, as DAA objectives are not strictly convex, this means that many optima of eq. 4 can place a high weight on OOD responses.
We can demonstrate this using the regression interpretation from Hejna et al. [23]. Consider re-writing the DAA objective from eq. 4 in terms of preference query vectors $q$ which select the win response pair $(x,y^{w})$ and the loss response pair $(x,y^{l})$ from the prompt-response space. Each vector $q$ represents both the preferred and dis-preferred responses, with the entree corresponding to $(x,y^{w})$ being +1 and the entree corresponding to $(x,y^{l})$ being -1. We can then write the generalized DAA loss function with finite preference data as
@@
$\mathcal{L}_{\text{DAA}}(\pi_{\theta},\mathcal{D})=\sum_{i=1}^{|\mathcal{D}|}g\left(\beta q_{i}^{\top}(\log\pi(y|x)-\log\pi_{\text{ref}}(y|x))\right),\text{where }q_{i}[x,y]=\begin{cases}1\;\;\text{if }(x^{(i)},y^{(i)}_{w})=(x,y)\\ -1\>\text{if }(x^{(i)},y^{(i)}_{l})=(x,y)\\ 0\;\;\;\text{otherwise}\end{cases}$
@@
where the policy $\pi$ is a single vector of size $|\mathcal{X}\times\mathcal{Y}|$. In practice, the distributional constraint of $\pi$ also applies. Choosing $g$ to be the negative log sigmoid above recovers DPO with finite preferences, but also logistic regression with a data matrix $Q$ of shape $|\mathcal{D}|$ by $|\mathcal{X}\times\mathcal{Y}|$ constructed by stacking the aforementioned query vectors $q$. As $|\mathcal{X}\times\mathcal{Y}|>>|\mathcal{D}|$, this matrix is likely to have a non-trivial null space, making the problem not strictly convex. Thus, there are many possible policies $\pi$ that can achieve the same optima, some of which can place a high weight on out-of-distribution responses [23, 69]. For more details and constructions, we defer to Hejna et al. [23].

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has been crucial to the recent success of Large Language Models (LLMs), however, it is often a complex and brittle process. In the classical RLHF framework, a reward model is first trained to represent human preferences, which is in turn used by an online reinforcement learning (RL) algorithm to optimize the LLM. A prominent issue with such methods is reward over-optimization or reward hacking, where performance as measured by the learned proxy reward model increases, but true quality plateaus or even deteriorates. Direct Alignment Algorithms (DAAs) like Direct Preference Optimization have emerged as alternatives to the classical RLHF pipeline by circumventing the reward modeling phase. However, although DAAs do not use a separate proxy reward model, they still commonly deteriorate from over-optimization. While the so-called reward hacking phenomenon is not well-defined for DAAs, we still uncover similar trends: at higher KL budgets, DAA algorithms exhibit similar degradation patterns to their classic RLHF counterparts. In particular, we find that DAA methods deteriorate not only across a wide range of KL budgets but also often before even a single epoch of the dataset is completed. Through extensive empirical experimentation, this work formulates and formalizes the reward over-optimization or hacking problem for DAAs and explores its consequences across objectives, training regimes, and model scales.
@@ 3 Empirical Analysis of Overoptimization in DAAs
First, we examine the over-optimization problem in DAAs and compare it to those observed in traditional RLHF methods. All our experiments are carried out using the Reddit TL;DR summarization dataset [55] and the Pythia family of Large Language Models [6]. Additional plots illustrating similar over-optimization trends for Direct Alignment Algorithms on the Gemma2-2b model [61] and the Anthropic Helpfulness-Harmlessness dataset [5] are provided in Appendix F
@@ 4 Reward Exploitation in Direct Alignment Algorithms
While the phenomena observed in the previous section echo those observed in classical RLHF, their underlying causes may be distinct. Reward over-optimization in classical RLHF is largely attributed to querying a proxy reward function that is potentially OOD, while DAAs do not train a separate reward model. Instead, DAAs are generally understood as fitting an “implicit” reward model to preference data with the parameterization $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}{(y|x)}}$ using the objective in eq. 2. Therefore, the OOD behavior of the policy is inextricably linked to the OOD behavior of the implicit reward model. We demonstrate below that the reward modeling objective used is heavily under-constrained, allowing for a potentially large number of solutions that can place weight on OOD responses. This is especially problematic for DAAs which deterministically map the optimal policy from the “implicit” reward.
@@
Rank Deficiency with Finite Preferences. In DAAs, the language modeling problem is treated as a contextual bandit. However, the space of possible prompts $x\in\mathcal{X}$ and answers $y\in\mathcal{Y}$ are both exponentially large in sequence length. However, as highlighted by Tang et al. [59], DAAs often assume full support of the reference distribution when mapping from the implicit reward to the optimal policy $\pi$ by eq. 10. However, in practice such coverage is impossible. Instead, preference datasets cover a minuscule portion of the prompt-response space. Unfortunately, as DAA objectives are not strictly convex, their loss functions (eq. 4) can have multiple global optimas, which may be undesirable. We demonstrate this below, using the regression interpretation from Hejna et al. [23].
First, we re-write the DAA objective from eq. 4 using vectors in the prompt-response space $\mathcal{X}\times\mathcal{Y}$. Each preference query in the comparison dataset can be written as difference between indicator vectors, specifically $q_{i}=\mathds{1}\{(x,y)=(x^{(i)},y^{(i)}_{w})\}-\mathds{1}\{(x,y)=(x^{(i)},y^{(i)}_{l})\}$. This “query” vector simply selects the comparison from the prompt response space, with the entree corresponding to $(x,y^{w})$ being +1 and the entree corresponding to $(x,y^{l})$ being -1. Similarly, we can consider the learned policy to be a vector $\log\pi-\log\pi_{\text{ref}}\in\mathcal{X}\times\mathcal{Y}$, to which the distributional constraint also applies in practice. Our generalized DAA loss function can then be re-written as
@@
$\mathcal{L}_{\text{DAA}}(\pi_{\theta},\mathcal{D})=\sum_{i=1}^{|\mathcal{D}|}g\left(\beta q_{i}^{\top}\left(\log\pi(y|x)-\log\pi_{\text{ref}}(y|x)\right)\right),\text{ where }q_{i}[x,y]=\begin{cases}1\;\;\text{if }(x,y)=(x^{(i)},y^{(i)}_{w})\\ -1\>\text{if }(x,y)=(x^{(i)},y^{(i)}_{l})\\ 0\;\;\;\text{otherwise}\end{cases}$
@@
with finite data. Choosing $g$ to be the negative log sigmoid above recovers DPO with finite preferences, but also logistic regression with a data matrix $Q$ of shape $|\mathcal{D}|$ by $|\mathcal{X}\times\mathcal{Y}|$ constructed by stacking the aforementioned query vectors $q$. As $|\mathcal{X}\times\mathcal{Y}|>>|\mathcal{D}|$, this matrix is likely to have a non-trivial null space, making the problem not strictly convex. Thus, there are many possible policies $\pi$ that can achieve the same optima, some of which will place a high weight on out-of-distribution responses due to the distributional constraint of policy [23, 70]. To demonstrate this, we formalize the construction below.
Proposition 1
(Adapted from Hejna et al. [23]) Let $S$ be the set of win-or-lose prompt-response vectors $(x,y)$ in $\mathcal{D}$. Provided:
  1. 1.
    The intersection of the null space of $Q$, $N(Q)$, and the span of $S$, $\text{span}(S)$, is non-trivial.
  2. 2.
    For every $x$ there exists a response $y_{\text{OOD}}\in\mathcal{Y}$ that is not in the data, $(x,y_{\text{OOD}})\neq S$.
Then, there are infinite number of minima to eq. 4 which place weight on out-of-distribution responses $y$.
Proof. Let $\hat{\pi}$ be the minima of the DAA loss function. Choose a vector $u$ such that $u\in N(Q)$, $u\in\text{span}(S)$, and $u$ has at least one negative component. Modifying the log policy vector as $\log\hat{\pi}+u$ will not affect the DAA loss, as $u$ is in the null space of $Q$, but the log-probability of the policy will decrease for least one prompt-response pair in $S$ by construction. However, $e^{\log\hat{\pi}+u}$ may not integrate to one. To fix this, we can construct a second vector $v\in N(Q)$ using the $y_{\text{OOD}}$ at each $x$ such that $e^{\log\hat{\pi}+u+v}$ integrates to one. For more details, we refer the reader to Hejna et al. [23] Appendix A.3.
The second constraint of proposition 1 is often trivially satisfied by the dimension of the response space as we are unlikely to see every response to a prompt. The first constraint is harder, but can be satisfied by conflicting preferences. A trivial example which satisfies these constraints is a simple MDP in which there is only a single state (or prompt $x$), but three possible actions (or responses), $y_{1}$, $y_{2}$, and $y_{3}$. If we construct the preference dataset $\mathcal{D}=\{(y_{1}\succ y_{2}),(y_{2}\succ y_{1})\}$, omitting $x$ for brevity, then we satisfy the above conditions: the null space of $Q$ is non trivial in span of $y_{1}$ and $y_{2}$ and there is an out-of-distribution action $y_{3}$. In this setting, the DPO loss is minimized by both $\hat{\pi}(y|x)=(0.5,0.5,0)$ and $\hat{\pi}(y|x)=(0.0,0.0,1.0)$. In fact, it is minimized by infinitely many policies which place equal weight on $y_{1}$ and $y_{2}$. To demonstrate this effect in higher dimensions across a number of different DAA methods, we conduct experiments in a Toy MDP which bears resemblance to the language modeling setting.
@@ 6 Conclusion
Acknowledgments
This work has taken place in part in the Safe, Correct, and Aligned Learning and Robotics Lab (SCALAR) at The University of Massachusetts Amherst. SCALAR research is supported in part by the NSF (IIS-2323384), AFOSR (FA9550-20-1-0077), and the Center for AI Safety (CAIS). This work has taken place in part in the Rewarding Lab at UT Austin. The Rewarding Lab is supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), EA Ventures, Bosch, UT Austin’s Good Systems grand challenge, and Open Philanthropy.
@@ E.1 Designing a toy LLM MDP
Appendix F Overoptimization Trends in the Gemma2-2b Model and Anthropic-HH Dataset
We present KL divergence versus GPT-4 win rate plots in Figure 19 to illustrate overoptimization trends in Direct Alignment Algorithms for the Gemma2-2b model [61] and the Anthropic-HH dataset [5]. Results are shown for the DPO and SLiC variants, which sufficiently demonstrate that the overoptimization trends observed with the Pythia models are not specific to a single model or dataset. The figure illustrates the trade-off between KL divergence and GPT-4 win rate across different values of beta in the alignment objective.
Figure 19: KL divergence versus GPT-4 win rate for the Gemma2-2b model on the Anthropic-HH dataset. The left plot shows DPO results, and the right plot shows SLiC results.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper faithfully adheres to the claims and motivation in the abstract and provides proof and detailed empirical studies in support.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: A discussion of our limitations can be found as a separate section at the beginning of the appendix.
  9. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  10. 3.
    Theory Assumptions and Proofs
  11. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  12. Answer: [Yes]
  13. Justification: We provide proofs and empirical evidence to support all our theoretical results.
  14. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  15. 4.
    Experimental Result Reproducibility
  16. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  17. Answer: [Yes]
  18. Justification: We provide detailed guidance on reproducibility by specifying all datasets, code, and hyperparameters used in this work.
  19. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  20. 5.
    Open access to data and code
  21. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  22. Answer: [Yes]
  23. Justification: We have only used open-source models with open-source datasets for all aspects of the work. Please refer to section B for details on reproducing the results.
  24. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: We list detailed information about the training and test details in Section LABEL:. Our experiments use open-source datasets and models.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [No]
  33. Justification: Training Large Language models is time-consuming and compute-intensive. Our experiments do not run multiple seeds on one configuration due to limited computing and financial budget. Instead, the focus of this work is extensive evaluation across multiple configurations which we spent all our compute resources into. Our evaluation protocol is similar to prior influential works in RLHF [46, 21].
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: We provide information on compute resources in the experimental details section B in the appendix.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: We abide by the code of ethics in every respect.
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: We discuss societal impacts in Section A of the appendix.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: We use public models that are fine-tuned for alignment on open-source datasets. Our models do not contribute any additional risk over the base models as we are explicitly training for alignment.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: The pretrained models in this work come from the Pythia family all of which are classified under Apache License https://huggingface.co/EleutherAI/pythia-2.8b/tree/main. The TL;DR comparison dataset used in this work uses a modified MIT License https://github.com/openai/summarize-from-feedback/blob/master/LICENSE.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: We use open-source pretrained models and provide details to reproduce our fine-tuning experiments.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: We do not use crowdsourcing or research with human subjects in this work
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: We do not use crowdsourcing or research with human subjects in this work.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.